### v2

@@ 5.3 The finite-dimensional learning problem and its fidelity
$\|\big{(}\mathcal{P}_{Q}(a_{*}),\mathcal{P}_{Q}(b_{*})\big{)}\|^{2}_{\mathcal{H}_{m}}\leq C(a_{*},b_{*}),$
@@
Let the coefficients $(\widehat{a}_{Q},\widehat{b}_{Q})\in\mathcal{H}^{+,Q}_{m,R_{*}}$ be the unique solution to LPQ with ${\color[rgb]{0,0,0}\lambda=\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{2}}$ and $Q\in\mathbb{N}$ so that ${\color[rgb]{0,0,0}h_{Q}=\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{\frac{1}{2m-1}}}\leq 1$. With probability at least $1-\delta$, it holds that:
@@
where the constant $C(a_{*},b_{*})>0$ depends on $a_{*}$ and $b_{*}$ uniquely.
We conclude this section with a result which summarizes our essential contributions, showing how our learning approach may be leveraged for efficient observation/regulation of stochastic differential equations. First, we recall the setting introduced at the end of Section 3. Specifically, for every coefficients $(a,b)\in\mathcal{H}^{+}_{m}$, we denote by $X^{a,b}$ and $p_{a,b}$ respectively the (unique) solutions to SDE and SFPE with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$. By combining Theorem 5.19 with Corollary 3.3 (in particular, compare with (10); more details about this remark are provided in Section 8), we readily obtain the following summarizing result:
@@
Let $m\in\mathbb{N}$, $\alpha>0$, $R_{*}>0$, $p_{0}\in H^{2m+1}(\mathbb{R}^{n},\mathbb{R})$, and $(a_{*},b_{*})\in\mathcal{H}^{+}_{m,R_{*}}$ satisfy Assumptions $(A)$ and $(B)$, and denote by $X$ and $p$ respectively the (unique) solutions to SDE and SFPE with coefficients $(a_{*},b_{*})\in\mathcal{H}^{+}_{m,R_{*}}$. There exists a constant $C(a_{*},b_{*})>0$ which only depends on $a_{*}$ and $b_{*}$, such that by choosing the following learning parameters for the fixed precision parameters $10<\varepsilon,\delta<$ 1:
@@
  • ${\color[rgb]{0,0,0}\lambda=\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{2}}$, and $Q\in\mathbb{N}$ so that ${\color[rgb]{0,0,0}h_{Q}=\min\left(1,\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{\frac{1}{2m-1}}\right)}$,
@@
where $(\widehat{a}_{Q},\widehat{b}_{Q})\in\mathcal{H}^{+,Q}_{m,R_{*}}$ is the unique solution to the finite-dimensional learning problem LPQ, where the learning parameters have been selected as above. Therefore, for every $f\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ the following observation/regulation estimate holds:
@@ 6 Computational Considerations
Problem reformulation. Recall that we are naturally, though implicitly extending the mapping $\Phi_{Q}$ to zero outside $D$, smoothly. Thanks to the characterization form Eq. 21, which is valid for every $(a,b)\in\mathcal{H}^{+,Q}_{m,R_{*}}$, one easily shows that
$\mathcal{L}^{a,b}_{t}\widehat{p}(t,y)={\operatorname{tr}}(V(t,y)\,A)-U(t,y)B+r(t,y),\quad(t,y)\in[0,T]\times\mathbb{R}^{n},$
@@
$r(t,y)\triangleq\frac{\alpha}{2}\Delta\widehat{p}(t,y),$
@@
$U_{\ell}(t,y)\triangleq\sum_{i=1}^{n}\frac{\partial}{\partial x_{i}}\left[\widehat{p}(t,y)e^{\top}_{i}K_{D}((t,y),(t_{\ell},x_{\ell}))\right]\in\mathbb{R}^{n},\quad\ell,\ell^{\prime}=1,\dots,Q,$
$V_{\ell,\ell^{\prime}}(t,x)\triangleq\frac{1}{2}\sum_{i,j=1}^{n}\frac{\partial^{2}}{\partial y_{j}\partial y_{i}}\big{[}\widehat{p}(t,y)K_{D}((t,y),(t_{\ell},x_{\ell}))\times$
×$\hskip 193.74939pt\times K_{D}((t,y),(t_{\ell^{\prime}},x_{\ell^{\prime}}))e_{i}e^{\top}_{j}\big{]}\in\mathbb{R}^{n\times n}.$
Therefore, by denoting $q\triangleq\frac{\partial\widehat{p}}{\partial t}$, the cost in LPQ thus writes
@@
$=\sum_{i,i^{\prime},\ell,\ell^{\prime}=1}^{M}\sum_{j,j^{\prime}=1}^{N}\frac{\alpha_{i\ell}\alpha_{i^{\prime}\ell^{\prime}}}{N^{2}}\int_{0}^{T}\dot{k}_{m}(t-t_{i})\dot{k}_{m}(t-t_{i^{\prime}})\;\mathrm{d}t\ \times$
@@
Above we used the definition of $\widehat{p}$, and the fact that $c_{\ell}$ are minimum-norm kernel interpolators [48], i.e., they are given by $c_{\ell}(t)=\sum_{i=1}^{N}\alpha_{i,\ell}k_{m}(t-t_{i})$, where $\alpha_{i,\ell}=(K^{-1})_{i,\ell}$, whereas $K\in\mathbb{R}^{N\times N}$ is given by $K_{i,i^{\prime}}=k_{m}(t_{i},t_{i^{\prime}})$. By denoting
$R_{\ell,\ell^{\prime}}\triangleq\int_{0}^{T}\dot{k}_{m}(t-t_{i})\dot{k}_{m}(t-t_{i^{\prime}})\;\mathrm{d}t,\quad S_{\ell,j,\ell^{\prime},j^{\prime}}\triangleq\int_{\mathbb{R}^{n}}\rho_{R}(y-x_{\ell,j})\rho_{R}(y-x_{\ell^{\prime},j^{\prime}})\;\mathrm{d}y,$
@@
$\int_{[0,T]\times\mathbb{R}^{n}}q(t,y)^{2}\;\mathrm{d}t\;\mathrm{d}y=\sum_{\ell,\ell^{\prime}=1}^{M}\sum_{j,j^{\prime}=1}^{N}(K^{-1}RK^{-1})_{\ell,\ell^{\prime}}S_{\ell,j,\ell^{\prime},j^{\prime}}.$
@@
Considerations on the computational complexity. Since to achieve a desired value for the fill distance $Q=O(h_{Q}^{-(n+1)})$ number of centers are required, taken, e.g. on a grid of stepsize $h_{Q}$ [40], we thus need $Q=O\left(\varepsilon^{-\frac{n+1}{m-(n-1)/4}}\right)$ to meet the rates of Theorem 5.19. In particular, by dimensioning $M$ and $N$ according to Theorem 5.19 to achieve the desired precision $\varepsilon$, it is required that $Q^{4}\leq M^{2}N^{2}$. Therefore, the proposed algorithm enjoy the following performance:
@@ 8.1 The Fokker-Planck equation
We recall we fixed $m\in\mathbb{N}$ and a constant $\alpha>0$, and since we worked with diffusion coefficients which are never trivial, for every $(a,b)\in\mathcal{H}^{+}_{m}$, the coefficient $a$ has been replaced with the mapping
$a+\alpha I:[0,T]\times\mathbb{R}^{n}\to\textnormal{Sym}_{++}(n)\quad\textnormal{with}\quad a+\alpha I=\sqrt{a+\alpha I}\sqrt{a+\alpha I}.$
@@
A measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ solves the Stochastic Differential Equation with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$ if each process $X_{x}(t,\omega)\triangleq X(x,\omega)(t)$ is $\mathcal{F}$–progressively measurable for every $x\in\mathbb{R}^{n}$, and
$\textnormal{SDE}_{x}\quad\begin{cases}\mathrm{d}X_{x}(t)=b(t,X_{x}(t))\;\mathrm{d}t+\sqrt{(a+\alpha I)(t,X_{x}(t))}\;\mathrm{d}W_{t},\\[5.0pt] \mathbb{P}\big{(}X_{x}(0)=x\big{)}=1,\end{cases}$
holds in $(\Omega,\mathcal{G},\mathcal{F},\mathbb{P})$ for $\mu_{0}$-almost every $x\in\mathbb{R}^{n}$. A solution $X$ to SDE is unique if, for every measurable mapping $Y:\mathbb{R}^{n}\times\Omega\to S$ which solves SDE with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$, it holds that $X(x,\cdot)=Y(x,\cdot)$ a.s., for $\mu_{0}$-almost every $x\in\mathbb{R}^{n}$.
@@
For every $(a,b)\in\mathcal{H}^{+}_{m}$, there exists a unique measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ which solves SDE with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$. In addition, the following properties hold true for the mapping $X$:
@@
(24) $\mathbb{E}\Big{[}\mathfrak{d}(X_{x_{1}},X_{x_{2}})^{2}\Big{]}\leq C\big{(}\|(a,b)\|_{\mathcal{H}^{+}_{m}}\big{)}\|x_{1}-x_{2}\|^{2},$
where the constant $C\big{(}\|(a,b)\|_{\mathcal{H}^{+}_{m}}\big{)}>0$ depends on $(a,b)\in\mathcal{H}^{+}_{m}$ uniquely (see, e.g., [22]) [^2] Here, we use the fact that the mapping $y\mapsto\sqrt{a(t,y)+\alpha I}$ is Lipschitz, for every $t\in[0,T]$. This is a straightforward consequence of Theorem 2.1 and the fact that the mapping $A\in\textnormal{Sym}_{++}(n)\mapsto\sqrt{A}$ is Lipschitz on $\{A\in\textnormal{Sym}_{++}(n):\ \|A\|\leq\|a\|_{L^{\infty}}+\alpha,\ y^{\top}Ay\geq\alpha\|y\|^{2},y\in\mathbb{R}^{n}\}$. Indeed, the latter set is compact and the mapping $A\in\textnormal{Sym}_{++}(n)\mapsto\sqrt{A}$ is continuously differentiable.. Therefore, we define the mapping $X:\mathbb{R}^{n}\times\Omega\to S$ by $X(x,\omega)(t)\triangleq X_{x}(t,\omega)$. From the continuity of the sample paths of each $X_{x}$, one may show that
@@
$\mathbb{E}\left[\underset{t\in[0,T]}{\sup}\ \left\|\tilde{X}_{x}(t)-x-\int^{t}_{0}b(s,\tilde{X}_{x}(s))\;\mathrm{d}r-\int^{t}_{0}\sqrt{(a+\alpha I)(s,\tilde{X}_{x}(s))}\;\mathrm{d}W_{s}\right\|^{2}\right]\leq$
@@
$\hskip 75.34698pt+C\mathbb{E}\left[\int^{T}_{0}\left\|\sqrt{(a+\alpha I)(t,\tilde{X}_{x}(t))}-\sqrt{(a+\alpha I)(t,X_{x}(t))}\right\|^{2}\;\mathrm{d}t\right]$
@@
for some (overloaded) constant $C>0$, and therefore by combining (24) with (25) yields that each process $\tilde{X}_{x}$ satisfies SDEx, $x\in\mathbb{R}^{n}$. In turn, we showed the existence of a measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ which solves SDE with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$. By leveraging the same argument (and a routine application of Gronwäll’s inequality), one also shows the uniqueness of this mapping as in Definition 8.1.
@@
At this step, for every $(a,b)\in\mathcal{H}^{+}_{m}$, we denote the Kolmogorov generator by
$\mathcal{L}^{a,b}_{t}\varphi(y)\triangleq\frac{1}{2}\sum^{n}_{i,j=1}(a+\alpha I)_{ij}(t,y)\frac{\partial^{2}\varphi}{\partial y_{i}\partial y_{j}}(y)+\sum^{n}_{i=1}b_{i}(t,y)\frac{\partial\varphi}{\partial y_{i}}(y),\quad\varphi\in C^{2}(\mathbb{R}^{n},\mathbb{R}).$
Fix $(a,b)\in\mathcal{H}^{+}_{m}$, and assume we are given a measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ which solves SDE with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$. For $\varphi\in C^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R})$, a straightforward application of Itô’s formula to SDEx yields
@@
A narrowly continuous curve $\mu:[0,T]\to\mathcal{P}(\mathbb{R}^{n})$ is said to solve the Fokker-Planck Equation with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$ if
@@
Importantly, thanks to the regularity of the coefficients $(a,b)\in\mathcal{H}^{+}_{m}$, FPE can have one narrowly continuous solution $\mu$ at most, as we state in the following:
@@
Given any tuple of coefficients $(a,b)\in\mathcal{H}^{+}_{m}$, at most one narrowly continuous curve $\mu:[0,T]\to\mathcal{P}(\mathbb{R}^{n})$ can solve FPE with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$.
Our previous computations show that solutions to FPE may be obtained from solutions to SDE, and we now establish this process may be inverted. Specifically, motivated by the results in [43, 14], we prove that any narrowly continuous curve $\mu:[0,T]\to\mathcal{P}(\mathbb{R}^{n})$ solution to FPE with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$ is associated with a unique (in the sense of Definition 8.1) measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ which solves SDE with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$. We gather such result in the following theorem, which is a natural extension of [14, Theorem 2.6] to our setting:
@@
Let a narrowly continuous curve $\mu:[0,T]\to\mathcal{P}(\mathbb{R}^{n})$ be solution to FPE with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$ (which is unique thanks to Proposition 8.5). There exists a unique measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ which solves SDE with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$, and which satisfies the representation formula:
@@
Thanks to Theorem 8.2, we already know there exists a unique measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ which solves SDE with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$.

### v3

@@ 5.3 The finite-dimensional learning problem and its fidelity
${\color[rgb]{0,0,0}\|\big{(}\mathcal{P}_{Q}(a_{*})-\alpha I,\mathcal{P}_{Q}(b_{*})\big{)}\|^{2}_{\mathcal{H}_{m}}\leq C(a_{*},b_{*}),}$
@@
Let the coefficients $(\widehat{a}_{Q},\widehat{b}_{Q})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,Q}_{m,R_{*}}}$ be the unique solution to LPQ with ${\color[rgb]{0,0,0}\lambda=\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{2}}$ and $Q\in\mathbb{N}$ so that $h_{Q}=\left(\log\left(\frac{1}{\delta\varepsilon}\right)\varepsilon\right)^{\frac{2}{2m-1}}\leq 1$. With probability at least $1-\delta$, it holds that:
@@
where the constant $C(a_{*},b_{*})>0$ depends on $a_{*}$ and $b_{*}$ uniquely (in particular, $C(a_{*},b_{*})$ implicitly depends on the hyper-parameters $m$, $p_{0}$, and $\alpha$).
We conclude this section with a result which summarizes our essential contributions, showing how our learning approach may be leveraged for efficient observation/regulation of stochastic differential equations. First, we recall the setting introduced at the end of Section 3. Specifically, for every coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, we denote by $X^{a,b}$ and $p_{a,b}$ respectively the (unique) solutions to SDE and SFPE with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$. By combining Theorem 5.18 with Corollary 3.3 (in particular, compare with (10); more details about this remark are provided in Section 8), we readily obtain the following summarizing result:
@@
Let $m\in\mathbb{N}$, $\alpha>0$, $R_{*}>0$, $p_{0}\in H^{2m+1}(\mathbb{R}^{n},\mathbb{R})$, and $(a_{*},b_{*})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m,R_{*}}}$ satisfy Assumptions $(A)$ and $(B)$, and denote by $X$ and $p$ respectively the (unique) solutions to SDE and SFPE with coefficients $(a_{*},b_{*})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m,R_{*}}}$. There exists a constant $C(a_{*},b_{*})>0$ which only depends on $a_{*}$ and $b_{*}$, such that by choosing the following learning parameters for the fixed precision parameters $10<\varepsilon,\delta<$ 1:
@@
  • ${\color[rgb]{0,0,0}\lambda=\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{2}}$ and $Q\in\mathbb{N}$ so that $h_{Q}=\min\left(1,\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{\frac{1}{2m-1}}\right)$,
@@
where $(\widehat{a}_{Q},\widehat{b}_{Q})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,Q}_{m,R_{*}}}$ is the unique solution to the finite-dimensional learning problem LPQ, where the learning parameters have been selected as above. Therefore, for every $f\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ the following observation/regulation estimate holds:
@@ 6 Computational Considerations
Problem reformulation. Recall that we naturally extend the mapping $\Phi_{Q}$ in (21) to zero outside $D$, and that we arbitrarily fix a smooth cut-off function $\chi$ to define ${\color[rgb]{0,0,0}\mathcal{H}^{+,Q}_{m,R_{*}}}$, such as an easily numerically implementable mollifier, which amounts to an additional hyper-parameter. Thanks to the characterization form Eq. 21, which is valid for every $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,Q}_{m,R_{*}}}$, one easily shows that
${\color[rgb]{0,0,0}(\mathcal{L}^{a,b}_{t})^{*}\widehat{p}(t,y)={\operatorname{tr}}(V(t,y)\,A)-U(t,y)B+r(t,y),\quad(t,y)\in[0,T]\times\mathbb{R}^{n},}$
@@
${\color[rgb]{0,0,0}\displaystyle r(t,y)\triangleq\alpha\Delta\bigg{(}\bigg{(}1-\frac{\chi}{2}\bigg{)}\widehat{p}\bigg{)}(t,y),}$
@@
$U_{\ell}(t,y)\triangleq\sum_{i=1}^{n}\frac{\partial}{\partial y_{i}}\left[{\color[rgb]{0,0,0}\chi(y)}\widehat{p}(t,y)e^{\top}_{i}K_{D}((t,y),(t_{\ell},x_{\ell}))\right]\in\mathbb{R}^{n},\quad\ell,\ell^{\prime}=1,\dots,Q,$
$V_{\ell,\ell^{\prime}}(t,x)\triangleq\frac{1}{2}\sum_{i,j=1}^{n}\frac{\partial^{2}}{\partial y_{j}\partial y_{i}}\big{[}{\color[rgb]{0,0,0}\chi(y)}\widehat{p}(t,y)K_{D}((t,y),(t_{\ell},x_{\ell}))\times$
×$\hskip 193.74939pt\times K_{D}((t,y),(t_{\ell^{\prime}},x_{\ell^{\prime}}))e_{i}e^{\top}_{j}\big{]}\in\mathbb{R}^{n\times n},$
with $A\in\mathbb{R}^{Qn\times Qn}$, $A\succeq 0$, and $B\in\mathbb{R}^{Qn\times 1}$ our variables. Therefore, by denoting $q\triangleq\frac{\partial\widehat{p}}{\partial t}$, the cost in LPQ writes
@@
$=\sum_{i,i^{\prime},\ell,\ell^{\prime}=1}^{M}\sum_{j,j^{\prime}=1}^{N}\frac{\alpha_{i\ell}\alpha_{i^{\prime}\ell^{\prime}}}{N^{2}}\int_{0}^{T}{\color[rgb]{0,0,0}\dot{K}_{m+1}}(t-t_{i}){\color[rgb]{0,0,0}\dot{K}_{m+1}}(t-t_{i^{\prime}})\;\mathrm{d}t\ \times$
@@
Above we used the definition of $\widehat{p}$, and the fact that $c_{\ell}$ are minimum-norm kernel interpolators [48], i.e., they are given by $c_{\ell}(t)=\sum_{i=1}^{N}\alpha_{i,\ell}{\color[rgb]{0,0,0}K_{m+1}}(t-t_{i})$, where $\alpha_{i,\ell}=({\color[rgb]{0,0,0}G}^{-1})_{i,\ell}$, whereas ${\color[rgb]{0,0,0}G}\in\mathbb{R}^{N\times N}$ is given by ${\color[rgb]{0,0,0}G}_{i,i^{\prime}}={\color[rgb]{0,0,0}K_{m+1}}(t_{i},t_{i^{\prime}})$. By denoting
$R_{\ell,\ell^{\prime}}\triangleq\int_{0}^{T}{\color[rgb]{0,0,0}\dot{K}_{m+1}}(t-t_{i}){\color[rgb]{0,0,0}\dot{K}_{m+1}}(t-t_{i^{\prime}})\;\mathrm{d}t,\quad S_{\ell,j,\ell^{\prime},j^{\prime}}\triangleq\int_{\mathbb{R}^{n}}\rho_{R}(y-x_{\ell,j})\rho_{R}(y-x_{\ell^{\prime},j^{\prime}})\;\mathrm{d}y,$
@@
$\int_{[0,T]\times\mathbb{R}^{n}}q(t,y)^{2}\;\mathrm{d}t\;\mathrm{d}y=\sum_{\ell,\ell^{\prime}=1}^{M}\sum_{j,j^{\prime}=1}^{N}({\color[rgb]{0,0,0}G}^{-1}R{\color[rgb]{0,0,0}G}^{-1})_{\ell,\ell^{\prime}}S_{\ell,j,\ell^{\prime},j^{\prime}}.$
@@
Considerations on the computational complexity. Since to achieve a desired value for the fill distance $Q=O(h_{Q}^{-(n+1)})$ number of centers are required, taken, e.g. on a grid of stepsize $h_{Q}$ [40], we thus need $Q=O\left(\varepsilon^{-\frac{n+1}{m-(n-1)/4}}\right)$ to meet the rates of Theorem 5.18. In particular, by dimensioning $M$ and $N$ according to Theorem 5.18 to achieve the desired precision $\varepsilon$, it is sufficient that $Q^{4}\leq M^{2}N^{2}$. Therefore, the proposed algorithm enjoy the following performance:
@@ 8.1 The Fokker-Planck equation
We recall we fixed $m\in\mathbb{N}$ and a constant $\alpha>0$, such that, since we worked with diffusion coefficients that are never trivial, we replaced $\mathcal{H}^{+}_{m}$ with
${\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}\triangleq(\alpha I,0)+\Big{\{}(a,b)\in\mathcal{H}^{+}_{m}:\ a(t,x)\succcurlyeq 0,\ (t,x)\in[0,T]\times\mathbb{R}^{n}\Big{\}}.}$
@@
A measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ solves the Stochastic Differential Equation with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ if each process $X_{x}(t,\omega)\triangleq X(x,\omega)(t)$ is $\mathcal{F}$–progressively measurable for every $x\in\mathbb{R}^{n}$, and
$\textnormal{SDE}_{x}\quad\begin{cases}\mathrm{d}X_{x}(t)=b(t,X_{x}(t))\;\mathrm{d}t+{\color[rgb]{0,0,0}\sqrt{a(t,X_{x}(t))}}\;\mathrm{d}W_{t},\\[5.0pt] \mathbb{P}\big{(}X_{x}(0)=x\big{)}=1,\end{cases}$
holds in $(\Omega,\mathcal{G},\mathcal{F},\mathbb{P})$ for $\mu_{0}$-almost every $x\in\mathbb{R}^{n}$. A solution $X$ to SDE is unique if, for every measurable mapping $Y:\mathbb{R}^{n}\times\Omega\to S$ which solves SDE with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, it holds that $X(x,\cdot)=Y(x,\cdot)$ a.s., for $\mu_{0}$-almost every $x\in\mathbb{R}^{n}$.
@@
For every $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, there exists a unique measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ which solves SDE with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$. In addition, the following properties hold true for the mapping $X$:
@@
(26) ${\color[rgb]{0,0,0}\mathbb{E}\Big{[}\mathfrak{d}(X_{x_{1}},X_{x_{2}})^{2}\Big{]}\leq C\big{(}\alpha,\|(a-\alpha I,b)\|_{\mathcal{H}_{m}}\big{)}\|x_{1}-x_{2}\|^{2},}$
where the constant $C\big{(}\alpha,\|(a-\alpha I,b)\|_{\mathcal{H}_{m}}\big{)}>0$ depends on $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ uniquely (see, e.g., [22]) [^2] Here, we use the fact that the mapping $y\mapsto\sqrt{a(t,y)+\alpha I}$ is Lipschitz, for every $t\in[0,T]$. This is a straightforward consequence of Theorem 2.1 and the fact that the mapping $A\in\textnormal{Sym}_{++}(n)\mapsto\sqrt{A}$ is Lipschitz on $\{A\in\textnormal{Sym}_{++}(n):\ \|A\|\leq\|a\|_{L^{\infty}}+\alpha,\ y^{\top}Ay\geq\alpha\|y\|^{2},y\in\mathbb{R}^{n}\}$. Indeed, the latter set is compact and the mapping $A\in\textnormal{Sym}_{++}(n)\mapsto\sqrt{A}$ is continuously differentiable.. Therefore, we define the mapping $X:\mathbb{R}^{n}\times\Omega\to S$ by $X(x,\omega)(t)\triangleq X_{x}(t,\omega)$. From the continuity of the sample paths of each $X_{x}$, one may show that
@@
$\mathbb{E}\left[\underset{t\in[0,T]}{\sup}\ \left\|\tilde{X}_{x}(t)-x-\int^{t}_{0}b(s,\tilde{X}_{x}(s))\;\mathrm{d}r-\int^{t}_{0}{\color[rgb]{0,0,0}\sqrt{a(s,\tilde{X}_{x}(s))}}\;\mathrm{d}W_{s}\right\|^{2}\right]\leq$
@@
$\hskip 75.34698pt+C\mathbb{E}\left[\int^{T}_{0}\left\|{\color[rgb]{0,0,0}\sqrt{a(t,\tilde{X}_{x}(t))}-\sqrt{a(t,X_{x}(t))}}\right\|^{2}\;\mathrm{d}t\right]$
@@
for some (overloaded) constant $C>0$, and therefore by combining (26) with (27) yields that each process $\tilde{X}_{x}$ satisfies SDEx, $x\in\mathbb{R}^{n}$. In turn, we showed the existence of a measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ which solves SDE with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$. By leveraging the same argument (and a routine application of Gronwäll’s inequality), one also shows the uniqueness of this mapping as in Definition 8.1.
@@
At this step, for every $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, we denote the Kolmogorov generator by
$\mathcal{L}^{a,b}_{t}\varphi(y)\triangleq\frac{1}{2}\sum^{n}_{i,j=1}{\color[rgb]{0,0,0}a}_{ij}(t,y)\frac{\partial^{2}\varphi}{\partial y_{i}\partial y_{j}}(y)+\sum^{n}_{i=1}b_{i}(t,y)\frac{\partial\varphi}{\partial y_{i}}(y),\quad\varphi\in C^{2}(\mathbb{R}^{n},\mathbb{R}).$
Fix $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, and assume we are given a measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ which solves SDE with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$. For $\varphi\in C^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R})$, a straightforward application of Itô’s formula to SDEx yields
@@
A narrowly continuous curve $\mu:[0,T]\to\mathcal{P}(\mathbb{R}^{n})$ is said to solve the Fokker-Planck Equation with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ if
@@
Importantly, thanks to the regularity of the coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, FPE can have one narrowly continuous solution $\mu$ at most, as we state in the following:
@@
Given any tuple of coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, at most one narrowly continuous curve $\mu:[0,T]\to\mathcal{P}(\mathbb{R}^{n})$ can solve FPE with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$.
Our previous computations show that solutions to FPE may be obtained from solutions to SDE, and we now establish this process may be inverted. Specifically, motivated by the results in [43, 14], we prove that any narrowly continuous curve $\mu:[0,T]\to\mathcal{P}(\mathbb{R}^{n})$ solution to FPE with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ is associated with a unique (in the sense of Definition 8.1) measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ which solves SDE with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$. We gather such result in the following theorem, which is a natural extension of [14, Theorem 2.6] to our setting:
@@
Let a narrowly continuous curve $\mu:[0,T]\to\mathcal{P}(\mathbb{R}^{n})$ be solution to FPE with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ (which is unique thanks to Proposition 8.5). There exists a unique measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ which solves SDE with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, and which satisfies the representation formula:
@@
Thanks to Theorem 8.2, we already know there exists a unique measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ which solves SDE with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$.